OpenStudy (leahhhmorgannn):

How can you use the method of completing the square to transform any quadratic equation in x into an equation of the form (x – p)^2 = q that has the same solutions? I'm not really sure what this is asking... none of it really makes sense to me. I just need it cleared up and explained in simpler terms. @dan815

OpenStudy (dan815):

oaky so say you have aany quadratic equation they are all in the form ax^2+bx+c=y

OpenStudy (dan815):

if u wanna rewrite this in the form (x-p)^2=q then| ------------------------------ ax^2+bx+c=y x^2+b/ax+c/a =y/a x^2+Bx=C <--- lets just say i rearrange and call them these new variables >_> just for simplicity sake now.. x^2+Bx=C if you do (x+B/2)^2 since u know that we need an x^2 and Bx , but doing this gives extra terms so we simple add the extra term to the other side of the equation too to balanced it out Equation we started with : x^2+Bx=C (x+B/2)^2 = C+ B^2/4 another way to see this is so say we know the extra term we will get, lets just add it to both sides... x^2+Bx + B^2/4 =C +B^2/4 now factor the left side and you will see that x^2+Bx + B^2/4 = (x+B/2)^2

OpenStudy (leahhhmorgannn):

Okay, I get that. But if I had to answer the question in words only, how would I describe that process?

OpenStudy (dan815):

just read that whole thing completely, please!

OpenStudy (dan815):

it should be all you need

OpenStudy (dan815):

if aanthing in there doesnt make sense, ill be glad to clear it up, I kinda rushed so.. the grammar is pretty bad lol

OpenStudy (leahhhmorgannn):

Thank you :)